OpenStudy (anonymous):

Find the Volume of a cone with radius of 7 cm and a slant height of 25 cm.

OpenStudy (anonymous):

Well, you have the radius and the slant height, think of it as a right triangle and use the pythagorean theorem to find the height: So you get: \[r^2 + h^2= s^2 \rightarrow h = \sqrt{s^2 - r^2}\] Once you get height, volume of a cone is given by: \[V \frac{ 1 }{ 3}\pi *r^2*h\]

OpenStudy (anonymous):

so whats the answer?? @Zelda

OpenStudy (anonymous):

@faithelizabeth, I'm not going to do the work for you... I gave you what you needed. Now plug in and check yourself.

OpenStudy (anonymous):

274 is the answer hun

OpenStudy (anonymous):

for height anyways